remi2019:

Solve for x: 7x + 15 = 38 answer choices 23/7 53/7 23 53

remi2019:

plz help i am confused

supie:

\(7x+15=38\) \(-15\), both sides: \(7x+15−15=38−15=\) \(7x=23\) \(÷7\) (using fraction bars [fraction is division]) \(\LARGE\frac{7x}{7}=\frac{23}{7}\) \(\LARGE x= \frac{?}{?}\)

supie:

You get it?

supie:

If you don't understand something lmk.

remi2019:

not really

supie:

Alright, what are you confused on?

remi2019:

the last part

supie:

The last part all we were doing is dividing by `7`

remi2019:

oh so it is 23/7

supie:

Fractions are division so we use the fraction bar then \(\frac{7x}{7}\) cancels. so then all we have left is \(\frac{23}{7}\)

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @remi2019 oh so it is 23/7 \(\color{#0cbb34}{\text{End of Quote}}\) Yes .